# Beyond the Carnot limit: work extraction via an entropy battery

### v1

@@ Appendix G Endoreversible efficiency for two thermal reservoirs
In this section we will show that a Carnot like heat engine running between two thermal reservoirs using the models developed in this paper satisfy the maximum power efficiency limit for a single endoreversible cycle. With the entropy battery being an extension of these models into the spin angular momentum degree of freedom, then the entropy battery represents the efficiency achievable at maximum power when acting reversibly.
@@
For this heat engine, the reservoirs have temperatures $\tau_{\text{env}},\tau_{\text{batt}}$ with $\tau_{\text{env}}>\tau_{\text{batt}}$. These reservoirs thermalize to a final temperature $\tau_{f}$ as the unitary heat engines run. The unitless bosonic heat (2) for $\tau\gg 0$ is
@@
to first order. The heat released by the hot reservoir at temperature $\tau_{\text{env}}$ is therefore
$Q_{\text{env}}=Q(\tau_{f})-Q(\tau_{\text{env}})=(d-1)\Delta\tau_{f,\text{env}},$
where $\Delta\tau_{i,j}=\tau_{i}-\tau_{j}$. Likewise, the heat absorbed by the cold reservoir at temperature $\tau_{\text{batt}}$ is
$Q_{\text{batt}}=Q(\tau_{f})-Q(\tau_{\text{batt}})=(d-1)\Delta\tau_{f,\text{batt}}$
@@
$W=Q_{\text{env}}+Q_{\text{batt}}=(d-1)(2\tau_{f}-\tau_{\text{env}}-\tau_{\text{batt}})$
@@
$\eta=\frac{W}{Q_{\text{env}}}=\frac{2\tau_{f}-\tau_{\text{env}}-\tau_{\text{batt}}}{\tau_{f}-\tau_{\text{env}}}$
For the endoreversible heat engine with no spin waste bath, the final temperature is a standard result $\tau_{f}=\sqrt{\tau_{1}\tau_{2}}$, which is calculated based on the conditions of reversibility $\Delta\mathbb{S}=0$ and eq.(2). See [41] for details of the derivation. Upon substitution, and with some algebraic manipulation, we obtain the efficiency
$\eta=1-\sqrt{\frac{\tau_{\text{batt}}}{\tau_{\text{env}}}}.$
which is the energy efficiency at maximum power [39, 41]. The results in Fig.3 uses an endoreversible cycle, even as the number of spin states increases. Therefore, the results in Fig.3 represent the efficiency at maximum power.

### v2

@@ Appendix G Endoreversible efficiency of the entropy battery
In this section we will show that a Carnot like heat engine running between two thermal reservoirs using the models developed in this article satisfy the maximum power efficiency limit for a single endoreversible cycle. With the entropy battery being an extension of these models into the spin angular momentum degree of freedom, then the entropy battery represents the efficiency achievable at maximum power when acting reversibly.
@@
For this scenario, the reservoirs have temperatures $\tau_{H},\tau_{C}$ with $\tau_{H}>\tau_{C}$. These reservoirs thermalize to a final temperature $\tau_{f}$ as the unitary heat engines run. The unitless bosonic heat (2) for $\tau\gg 0$ is
@@
to first order. The heat released by the hot reservoir at temperature $\tau_{H}$ is therefore
$Q_{H}=Q(\tau_{f})-Q(\tau_{H})=(d-1)\Delta\tau_{f,H},$
where $\Delta\tau_{i,j}=\tau_{i}-\tau_{j}$. Likewise, the heat absorbed by the cold reservoir at temperature $\tau_{C}$ is
$Q_{C}=Q(\tau_{f})-Q(\tau_{C})=(d-1)\Delta\tau_{f,C}$
@@
$W=Q_{H}+Q_{C}=(d-1)(2\tau_{f}-\tau_{H}-\tau_{C})$
@@
$\eta=\frac{W}{Q_{H}}=\frac{2\tau_{f}-\tau_{H}-\tau_{C}}{\tau_{f}-\tau_{C}}$
For the endoreversible heat engine with no spin reservoir, the final temperature is a standard result $\tau_{f}=\sqrt{\tau_{H}\tau_{C}}$, which is calculated based on the conditions of reversibility $\Delta\mathbb{S}=0$ and eq.(2). See [51] for details of the derivation. Upon substitution, and with some algebraic manipulation, we obtain the efficiency
$\eta=1-\sqrt{\frac{\tau_{C}}{\tau_{H}}}.$
which is the energy efficiency at maximum power [52, 51].
Introducing the spin reservoir, the condition of reversibility becomes
$\Delta\mathbb{S}_{H}(\tau_{f})+\Delta\mathbb{S}_{C}(\tau_{f})+\Delta\mathbb{S}_{s}(\tau_{f})=0,$ (62)
where we can use the analytic bosonic entropy equation (25) to re-express the changes in entropy
$\Delta\mathbb{S}_{i}(\tau_{f})$ $=\sum_{j=1}^{d_{i}-1}\left(\frac{j}{2\tau_{f}}\coth(j/2\tau_{f})-\frac{j}{2\tau_{i}}\coth(j/2\tau_{i})\right$. (63)
$\left.+\ln\left(\frac{\sinh(j/2\tau_{i})}{\sinh(j/2\tau_{f})}\right)\right).$
This can be approximated to first order for $\tau_{i},\tau_{f}\gg 0$,
$\Delta\mathbb{S}_{i}(\tau_{f})\approx(d_{i}-1)\ln\left(\frac{\tau_{f}}{\tau_{i}}\right)$ (64)
giving an approximate final temperature
$\tau_{f}\approx\left(\tau_{H}^{d_{H}-1}\tau_{C}^{d_{C}-1}\tau_{s}^{d_{s}-1}\right)^{1/(d_{H}+d_{C}+d_{s}-3)},$ (65)
which is a weighted geometric mean of the initial temperatures. Setting $d_{i}$ all equal, then we find $\tau_{f}\approx\sqrt[3]{\tau_{H}\tau_{C}\tau_{s}}$, which is consistent with the final temperature of a three body system, each being the same size, yet unequal temperatures [51]. The implication of this result is that the energy efficiency obtained in §IV is for endoreversible engines running at maximum power. However this result does not determine the actual power of the engine. As such, we make no claim as to the actual power of the entropy battery, the analysis of which would require finite-time thermodynamics [69]. We therefore consider the entropy battery as operating at maximum power assuming it runs slowly and isolated enough to fulfill the condition of reversibility (62).